OpenStudy (abmon98):

2^x*2^(2x+1)=128 Solve equation.

OpenStudy (amoodarya):

\[2^x*2^{2x+1}=128\\2^x*2^{2x}*2^1=128\\2^x*2^{2x}=64\\2^{x+2x}=64\\2^{3x}=2^6\\3x=6\\x=2\]

OpenStudy (abmon98):

Do you how to solve the problem using logarithms and thank you for your help

OpenStudy (amoodarya):

do you wanna solve this by logarithm?

OpenStudy (abmon98):

yes

OpenStudy (amoodarya):

\[2^{x}*2^{2x+1}=128\\ \log_{2}\\ \log_{2}2^{x+2x+1}=\log_{2}128\\log_{2}2^{3x+1}=\log_{2}2^7 \]

OpenStudy (amoodarya):

so now 3x+1=7

OpenStudy (amoodarya):

\[(3x+1)\log_{2}2=7\log_{2}2\\log_{2}2=1\\3x+1=7\]

OpenStudy (abmon98):

Thank you so much @amoodarya